### v1

# MLM: Learning Multi-task Loco-Manipulation Whole-Body Control for Quadruped Robot with Arm

Abstract
Whole-body loco-manipulation for quadruped robots with arm remains a challenging problem, particularly in achieving multi-task control. To address this, we propose MLM, a reinforcement learning framework driven by both real-world and simulation data. It enables a six-DoF robotic arm–equipped quadruped robot to perform whole-body loco-manipulation for multiple tasks autonomously or under human teleoperation. To address the problem of balancing multiple tasks during the learning of loco-manipulation, we introduce a trajectory library with an adaptive, curriculum-based sampling mechanism. This approach allows the policy to efficiently leverage real-world collected trajectories for learning multi-task loco-manipulation. To address deployment scenarios with only historical observations and to enhance the performance of policy execution across tasks with different spatial ranges, we propose a Trajectory-Velocity Prediction policy network. It predicts unobservable future trajectories and velocities. By leveraging extensive simulation data and curriculum-based rewards, our controller achieves whole-body behaviors in simulation and zero-shot transfer to real-world deployment. Ablation studies in simulation verify the necessity and effectiveness of our approach, while real-world experiments on the Go2 robot with an Airbot robotic arm demonstrate the policy’s good performance in multi-task execution.
Index Terms:
Reinforcement learning, legged robot, multi-task loco-manipulation, whole-body control.
I Introduction
Quadruped robot motion control, especially through reinforcement learning (RL), has gained significant attention and development, enhancing their diverse and impressive capabilities [1, 2, 3, 4, 5, 6]. A six-degree-of-freedom (DoF) robotic arm can offer more possibilities for enabling quadruped robots to perform manipulations. By integrating the locomotion of a quadruped robot with the manipulation capabilities of a robotic arm, a broader range of tasks can be accomplished compared to using a quadruped robot alone [7, 8].
However, achieving multiple loco-manipulation tasks with a single policy presents a challenge for quadruped robots. Employing a single policy across multiple tasks can reduce control complexity by eliminating the need for task-specific policies. However, balancing multiple tasks remains a problem that has not been adequately addressed by existing methods. Consequently, an adaptive balancing mechanism is essential. The method for achieving adaptive balance requires further research. In simulation-driven RL paradigms, the virtual assets and rewards specific to different tasks often vary, and careful, personalized design typically leads to a significant workload and lacks generalizability. Furthermore, the gap of visual information required for manipulation is often difficult to address due to significant differences in texture, color, lighting, and other factors between simulation and the real world [9]. A feasible solution is to distill multiple tasks into multiple trajectories learning. This avoids incorporating observational data of the task objectives (such as visual information) into the control, instead replacing it with trajectories.
Figure 1: The robot showcases its ability to perform multi-task whole-body loco-manipulation in the real world.
The acquisition of trajectories can be achieved through various methods, including UMI-like approaches [10, 11], teleoperation [12, 13], exoskeleton-based systems [14]. It is important to note that we do not collect whole-body trajectory data for quadruped loco-manipulation, which distinguishes our approach from imitation learning. This is due to the challenges of species selection for data collection and the inability to obtain such data from real robots. Thus, we use the UMI-like approach to obtain only the end-effector 6D manipulation trajectories, a lightweight solution independent of the robots.
Integrating real-world multi-task manipulation trajectories into simulation and leveraging the simulation to efficiently generate training data that interacts with the environment is a feasible approach. Furthermore, by abstracting multiple tasks into multiple trajectories and designing a unified task reward, it can guide the quadruped robot in generating coordinated whole-body behaviors under a single policy to accomplish the tasks. These multi-task data from the real world provide spatial trajectories that align with the dynamics of the physical environment and are distributed across the robot’s manipulation space, which is essential for learning whole-body behaviors and generalizing to unseen trajectories [15]. This contrasts with sampling random spatial curves in simulation, which lack real-world dynamics like velocity and acceleration.
It is important to note that in deployment schemes that only have access to historical trajectories, such as teleoperation, future trajectory information cannot be used as observation. However, future information is crucial for the stability of whole-body execution tasks, as it indicates the direction of the next step in the action [16]. Thus, a trajectory prediction mechanism is necessary.
To address the aforementioned challenges, we introduce MLM, a learning-based multi-task loco-manipulation method. We integrate Fast-UMI [11] to provide manipulation trajectories. In the simulation, we train the control policy and transfer it to the real world. Our contributions are summarized as follows:
  • We propose a Trajectory-Velocity Prediction-based framework for training a multi-task loco-manipulation controller, driven by both real-world and simulation data, with trajectory interface deployment enabled through teleoperation and a diffusion policy.
  • We introduce a multi-task real-world trajectory data library, where adaptive and curriculum-based sampling enables the robot to balance performance across tasks and learn multi-task manipualtion with a single policy network.
  • We deploy a real quadruped robot with arm and validate the excellent performance of our method in whole-body loco-manipulation control across multiple tasks.
II Related Work
II-A RL-based Legged Locomotion
The success of RL-based motion controllers for legged robots, compared to model-based approaches, stems from the following key advantages. RL trains control policies using large amounts of data collected from simulated robot-environment interactions, capturing the dynamics of these interactions while leveraging far more data than real-world training allows [1, 3, 17]. Furthermore, reward-driven policy networks provide an end-to-end mapping from states to joint actions, eliminating the need for costly optimization during inference [18, 4].
Building on the parallel dynamic simulator Isaac Gym, a quadruped robot was trained to walk within minutes [19]. An actuator network was employed to model the complex actuator dynamics of the ANYmal robot, thereby enabling sim-to-real transfer [20]. Extending this approach, privileged learning was introduced to train the ANYmal robot, successfully achieving a blind motion controller for complex terrains [1]. Additionally, a method named RMA was proposed to develop robust quadruped locomotion controllers [3]. To facilitate gait learning and improve adaptation to complex terrains, gait priors were introduced [21, 2, 22], allowing robots to leverage motion data from flat terrains and rapidly converge to the desired gait pattern. While these efforts have significantly advanced the locomotion capabilities of legged robots [23, 24, 25, 26], the development of their loco-manipulation abilities still requires further exploration.
II-B Legged Loco-Manipulation
To endow the manipulation capabilities of legged robots, researchers have integrated the mobility of legged platforms with the manipulation abilities of robotic arms. For instance, in [27], footstep planning was addressed, and a hierarchical optimization-based controller was developed to enable a quadruped robot with a six-DoF robotic arm to perform manipulation. The study in [7] proposed a approach that searches for whole-body trajectories and determines contact schedules to solve mobile manipulation tasks within predefined environments. Nonetheless, these optimization-based approaches often require task-specific design, presenting significant challenges in integrating the execution of multiple tasks within a single control policy. Additionally, they incur substantial computational overhead due to the optimization processes involved.
In recent years, several studies have focused on employing RL to achieve end-to-end control of legged robots with robotic arms and exploring the execution of policies across multiple tasks. In [28], the arm’s motion is modeled as predictable external torques that the RL-trained locomotion policy uses to compensate for the manipulator’s influence. Nevertheless, this non-whole-body approach limits the ability of the arm to utilize the mobility and posture adjustment capabilities of the quadruped robot to expand its manipulation range. To achieve whole-body control, an RL-based Regularized Online Adaptation method was proposed, successfully achieving sim-to-real transfer for quadruped loco-manipulation [29]. Despite this, the manipulation space defined under the arm base are susceptible to the influence of body shaking and are not stable enough. We utilize world space to address this issue, as the operational objectives are independent of body movement. In [30], a nonlinear reward fusion module was proposed to achieve whole-body behaviors for a wheeled quadrupedal manipulator. However, it did not integrate automatic execution of the task. In [31], an RL-based approach that integrates visual information for autonomous loco-manipulation execution was proposed. Yet, the inverse kinematics (IK) strategy is also susceptible to locomotion interference due to the arm base being affected by the body’s movement. In [16], a vision-driven, trajectory-based loco-manipulation in the task space was realized for a quadruped robot. However, it did not consider the integration of multiple tasks and the balance of performance across tasks within a single policy. Additionally, it is only applicable to trajectory generation methods that involve future trajectory reasoning, such as diffusion policy (DP). Teleoperation that only involve historical trajectories are not suitable.
III Whole-body Control Method
We develop a real-world and simulation dual data driven RL framework for quadruped robot whole-body loco-manipulation control. During training, a library containing real-world manipulation trajectories for multiple tasks is introduced, and an end-to-end Trajectory-Velocity Prediction policy network is trained for control. The pipeline is illustrated in Figure 2.
Figure 2: Overview of the training and inference pipeline. We use reinforcement learning to train a whole-body control policy. A trajectory library, together with an adaptive sampling curriculum, supplies a multiple set of real-world manipulation trajectories for the policy to track. A normalized autoencoder encodes historical target trajectories, implicitly conveying information about upcoming targets; alternatively, the policy can receive future trajectories directly, enabling two deployment modes: teleoperation and DP generation. We also train an estimator to estimate base linear velocity $\hat{v}_{t}^{p}$ using supervised learning.
III-A Whole-body Loco-Manipulation Policy
We eliminate the requirement for object and environment perception in simulation, which speeds up training and avoids the large sim-to-real gap caused by visual perception. Since information is partially observable, we model the loco-manipulation task as a partially observable Markov decision process (POMDP) [17, 22]. We employ RL to train a leg-arm whole-body policy $\pi_{\theta}$, aiming to find optimal policy parameters $\theta$ that maximize the discounted expected return $\mathbb{E}_{\mathbf{\tau}(\pi_{\theta})}[\sum_{t=0}^{T}{\gamma^{t}r_{t}}]$, where $r_{t}$ is the reward at time $t$, and $\gamma$ is the discount factor ($\gamma=0.99$).
We introduce an asymmetric Actor-Critic (AAC) framework to train the loco-manipulation policy. This approach is designed to address the POMDP problem while also avoiding the complexity of the two-stage training process [4].
State Space: The full state for the policy includes the quadruped robot’s proprioception $\boldsymbol{o}_{t}^{q}\in{\mathbb{R}^{42}}$, the six-DoF robotic arm’s proprioception $\boldsymbol{o}_{t}^{a}\in{\mathbb{R}^{18}}$, the trajectory pose sequence $o_{t-3:t+4}^{tcp}\in{\mathbb{R}^{72}}$, the privileged state $\boldsymbol{s}_{t}^{p}\in{\mathbb{R}^{21}}$, and the terrain elevation data $\boldsymbol{s}_{t}^{e}\in{\mathbb{R}^{187}}$.
Specifically, $\boldsymbol{o}_{t}^{q}$ includes the quadruped’s body angular velocity $\boldsymbol{\omega}_{t}^{q}\in{\mathbb{R}^{3}}$, the gravity vector $\boldsymbol{g}_{t}^{q}\in{\mathbb{R}^{3}}$ representing its orientation, leg joint positions $\boldsymbol{\theta}_{t}^{q}\in{\mathbb{R}^{12}}$ and velocities $\dot{\boldsymbol{\theta}}_{t}^{q}\in{\mathbb{R}^{12}}$, and the previous time step’s leg joint actions $\boldsymbol{a}_{t-1}^{q}\in{\mathbb{R}^{12}}$. The arm proprioception $\boldsymbol{o}_{t}^{a}$ contains the arm joint positions $\boldsymbol{\theta}_{t}^{a}\in{\mathbb{R}^{6}}$, velocities $\dot{\boldsymbol{\theta}}_{t}^{a}\in{\mathbb{R}^{6}}$, and last arm joint actions $\boldsymbol{a}_{t-1}^{a}\in{\mathbb{R}^{6}}$. The trajectory pose sequence $o_{t-3:t+4}^{tcp}$ comprises Tool Center Points (TCP) over past and subsequent steps relative to time $t$, each including a 3D position and the first six elements of the rotation matrix. The privileged state $\boldsymbol{s}_{t}^{p}$, covering both robot and environment, includes the quadruped’s body linear velocity $\boldsymbol{v}_{t}^{q}\in{\mathbb{R}^{3}}$, contact forces $\boldsymbol{f}^{c}_{t}\in{\mathbb{R}^{12}}$, external force $\boldsymbol{f}^{e}_{t}\in{\mathbb{R}^{3}}$, and its position $\boldsymbol{p}^{e}_{t}\in{\mathbb{R}^{3}}$ on the robot. Finally, $\boldsymbol{s}_{t}^{e}\in{\mathbb{R}^{187}}$ represents the heights of 187 terrain points evenly distributed around the robot’s coordinate frame.
Action Space: The whole-body policy outputs 12 joint position offsets for the legs, $\boldsymbol{a}_{t}^{q}\in{\mathbb{R}^{12}}$, and 6 joint position offsets for the arm, $\boldsymbol{a}_{t}^{a}\in{\mathbb{R}^{6}}$, both relative to their default positions. These offsets are added to the default positions and sent to the motors as target position commands for proportional-derivative (PD) control [30].
Reward: Maintaining coordinated loco-manipulation of quadruped robots is challenging. Building on [22], we use adversarial motion prior (AMP) to encourage quadruped gaits learning via style rewards. During policy training, an adversarial discriminator network is trained concurrently to provide style rewards, defined as $r^{s}=\max[0,1-0.25\times\left(d^{\rm score}-1\right)^{2}]$.
The primary reward of the policy, $r^{g}$, is the tracking of the robotic arm’s TCP position and orientation:
$r^{g}$ $=r^{g_{p}}\cdot r^{g_{o}}$ (1)
$r^{g_{p}}$ $=\exp({-(e_{1}^{p}/\sigma_{1}^{p}\oplus e_{2}^{p}/\sigma_{2}^{p}\oplus...\oplus e_{N}^{p}/\sigma_{N}^{p})})$
$r^{g_{o}}$ $=\exp({-(e_{1}^{o}/\sigma_{1}^{o}\oplus e_{2}^{o}/\sigma_{2}^{o}\oplus...\oplus e_{N}^{o}/\sigma_{N}^{o})})$
where $N$ is the total number of tasks. $\oplus$’ represents the concatenation operation. We have established a curriculum to update different tasks reward factor $\sigma_{n}^{p}$ and $\sigma_{n}^{o}$, more details are described in III-C. The constraint reward, $r^{l}$, is used to regulate the robot’s reasonable actions, energy limitations, and hardware constraints. All rewards are shown in Table I.
TABLE I: Rewards for learning whole-body loco-manipulation.
Term Equation Weight
$r^{g}$ (1) 2.0
$r^{l}$ $\|\boldsymbol{\tau}\|_{2}$ 1E41E-4
$\|\boldsymbol{\ddot{q}}\|_{2}$ 2.5E72.5E-7
$\|\boldsymbol{a}_{t-1}-\boldsymbol{a}_{t}\|_{2}$ 0.1
$n_{collision}$ 0.1
$\|\max\left(\left|\boldsymbol{\tau}\right|-\boldsymbol{\tau}^{limit},0\right)\|_{2}$ 0.05
$\|\max\left(\left|\boldsymbol{\dot{q}}\right|-\boldsymbol{\dot{q}}^{limit},0\right)\|_{2}$ 0.5
$r^{s}$ $\max[0,1-0.25\times(d^{\rm score}-1)^{2}]$ 0.5
Terrain Curriculum: In addition to the enhanced ability to adjust its posture with more degrees of freedom, the leg configuration of quadruped robots enables flexible adaptation to complex terrains. To enable a quadruped robot to perform loco-manipulation tasks in complex terrains, we introduce terrain curriculum. This approach allows the robot to progressively learn whole-body behaviors across several types of terrain, starting from easier to more difficult conditions.
These terrains include randomly undulating surfaces, slopes, discrete obstacles, stairs. Unlike pure locomotion tasks, we have removed the terrain traversal update mechanism, and instead, we use reward threshold updates. Specifically, we monitor the TCP tracking reward within an episode, and when it reaches the threshold, the robot progresses to the next level of the curriculum in the subsequent episode; otherwise, it regresses to the previous level. We have found that this mechanism effectively ensures the robot maintains a good tracking performance while preserving the stability of the base.
Domain Randomization: To reduce the sim-to-real gap, we randomize robot hardware, control, and environment parameters within specified ranges during training to ensure the simulation covers or approximates real-world conditions. As shown in Table II. This includes PD gains, link mass, payload mass and position, ground friction, and motor strength. Additionally, an observation delay of 0 to 0.02 seconds is introduced to mitigate the system’s non-real-time behavior.
TABLE II: Randomized parameters and their ranges.
Parameters Range [Min, Max] Unit
[…]
Payload Mass [0, 2] Kg
Payload Position [-0.05, 0.05] relative to base origin m
Ground Friction [0.05, 2.0] -
Motor Strength [0.8, 1.2] -
Observation Delay [0, 0.02] s
III-B Network
The training framework consists of a policy network and a Critic network. The policy network acts as the Actor and controller, receiving only observable states as input.
Trajectory-Velocity Prediction Policy Network: We propose a Trajectory-Velocity Prediction Policy Network (TVP), which predicts unobservable target trajectories and body velocity, integrates additional proprioception, and outputs joint actions. A trajectory encoder $E_{\theta_{1}}$ extracts features from the historical pose frames. Then, these features are fed into a decoder $D_{\theta_{2}}$ to predict the subsequent trajectory sequence, as shown in (2). This prediction is essential, particularly in scenarios such as teleoperation, where only historical trajectory records are available. Alternatively, when deploying the strategy that utilize DP, the future trajectory can be provided directly, and the TVP will be skipped.
$\hat{z}_{t}^{tcp}=E_{\theta_{1}}\left(o_{t-3:t}^{tcp}\right)$ $\hat{o}_{t+1:t+4}^{tcp}=D_{\theta_{2}}\left(\hat{z}_{t}^{tcp}\right)$ (2)
We introduce a Normalized Autoencoder (NAE) to formulate the structure and loss function for the trajectory predictor. The reconstruction loss measures how well the NAE reconstructs the future trajectory from the latent embedding $\hat{z}_{t}^{tcp}$:
$\mathcal{L}_{NAE}$ $=\mathcal{L}_{rec}=\mathrm{MSE}\left(\hat{o}_{t+1:t+4}^{tcp},o_{t+1:t+4}^{tcp}\right)$ (3)
Specifically, the MSE is used to minimize the discrepancy between the predicted future trajectory and the ground truth. Importantly, we apply the tanhtanh activation function to the encoded latent embedding for normalization, constraining them within a limited range. This helps prevent outliers that lie outside the expected distribution, thus aiding in stabilizing the training process. Additionally, we update the trajectory predictor, constructed by NAE, concurrently with other networks. On one hand, this facilitates accurate prediction of future trajectories; on the other hand, it generates useful latent embedding for use by the low-level network, thereby promoting the learning of the desired actions.
The memory encoder encodes the last five frames of proprioceptive observations. And an estimator is trained through supervised learning to estimate the body’s linear velocity:
$\hat{z}_{t}^{prop}=E_{\theta_{3}}\left(o_{t-4:t}\right)$ $\hat{v}_{t}^{p}=EST_{\theta_{4}}\left(\hat{z}_{t}^{prop}\right)$ (4)
$\mathcal{L}_{est}=\mathrm{MSE}\left(\hat{v}_{t}^{p},{v}_{t}^{p}\right)$
The estimation is important because loco-manipulation tasks involve both long-range movement and short-range local operations. The estimated linear velocity of the body effectively captures the distinction between these two types of tasks, enabling more accurate execution of these different loco-manipulation tasks. The encoded embedding, along with other proprioceptive data and previous actions, is fed into a low level network, which then outputs the mean vector $\boldsymbol{\mu}_{t}^{b}\in{\mathbb{R}^{18}}$ of a Gaussian distribution $\boldsymbol{a}_{t}^{b}\sim\mathcal{N}\left(\boldsymbol{\mu}_{t}^{b},\boldsymbol{\sigma}^{b}\right)$, where $\boldsymbol{\sigma}^{b}\in{\mathbb{R}^{18}}$ represents the variance of the action, optimized by proximal policy optimization (PPO).
Critic Network: The Critic network comprises a trajectory encoder $E_{\psi_{1}}$ and a height encoder $E_{\psi_{2}}$, which perform dimensionality reduction and representation. $\psi_{1}$ and $\psi_{2}$ represent the Critic network parameters. These encoded features, combined with proprioceptive observations, previous actions and privilege information, are used to compute the state value function $V_{t}$, assessing the expected reward for the current state:
$V_{t}=E_{\psi_{3}}\left(E_{\psi_{1}}(o_{t-3:t+4}^{tcp})\oplus E_{\psi_{2}}(s_{t}^{e})\oplus o_{t}\oplus a_{t-1}\oplus s_{t}^{p}\right)$ (5)
III-C Trajectory Generation and Sampling
Trajectory generation encompasses both the trajectories learned during training and those tracked during deployment. The following details the generation and sampling strategies for these trajectories.
Trajectory Library in Training: We incorporate a trajectory library in whole-body policy training, which contains real-world TCP pose data from multiple manipulation tasks. During training, the initial points of these trajectories are set as the end-effector poses following episodes reset, preventing low-quality samples in the early stages.
We select trajectories from several typical manipulation tasks for learning [11]. Each task includes 200 trajectories, such as ‘clean table’, ‘pour coke’, ‘open container’, ‘pick bear’, ‘pick cup’, and ‘unplug charger’. More trajectories can be added to expand multi-task learning. We also incorporate ‘pushing’ data from [16]. These tasks cover repetitive motions, pick-and-place actions, rotational opening and closing, and long-distance movements, making them representative of common real-world operational tasks. The data are stored as separate pickle files for training. Furthermore, we find that with the guidance of the TCP trajectory tracking reward $r^{g}$, the robot can generalize to tracking other unseen trajectories to some extent.
Trajectory Sampling Update: The complexity of task trajectories varies: some involve more changes in both position and orientation, while others involve fewer changes. If trajectories are sampled uniformly across all tasks, the policy may tend to focus on tracking easier trajectories, as they lead to higher rewards, while performing poorly on more challenging task. This can result in the policy getting trapped in local optima.
To address this, we propose an adaptive trajectory sampling and reward factor updating mechanism, which differs from the uniform sampling throughout the whole training. In the initial stage, trajectories from all tasks are sampled with equal probability. Meanwhile, the average tracking rewards for each task are recorded. When a task’s tracking reward changes, its sampling probability is updated accordingly as follows, prioritizing more challenging trajectories to enhance learning.
${P}^{\prime}_{n}=\frac{1}{1+\exp\left(k_{n}\left(\bar{r}_{n}^{g}-\lambda_{n}\right)\right)}$ (6)
$P_{n}=\frac{P^{\prime}_{n}}{{\textstyle\sum_{i=1}^{N}}P^{\prime}_{i}},i,n=1,.,N$
where $\bar{r}_{n}^{g}$ denotes the average tracking reward for each task trajectory. To prevent forgetting, this mechanism retains a minimum probability for each trajectory.
Meanwhile, each task trajectory adopts its own reward coefficient factor. Specifically, after repeated training and adjustments, a predefined set of average pose error thresholds and task reward factors is established as $\left\{\left(T_{k,n},\sigma_{k,n}\right)\right\}$. The average pose error for each task is calculated as follows, including Euclidean distance in displacement space and angular error in rotation space.
$e_{n}^{p}=\|\boldsymbol{p}_{n,{\rm tcp}}^{\rm target}-\boldsymbol{p}_{n,{\rm tcp}}\|_{2}$ (7)
$e_{n}^{o}={\arccos(0.5{\textstyle\sum_{i}}(\boldsymbol{\rm{R}}_{n,{\rm tcp}}^{\rm target}\cdot\boldsymbol{\rm{R}}_{n,{\rm tcp}}^{T})_{ii}-0.5)}$
where $\sigma_{n}^{p}$ and $\sigma_{n}^{o}$ are then updated according to (8).
[…]
The factor is adaptively updated based on changes in the reward values for each task trajectory. As the training progresses, this factor increases the tracking requirements for each task. The task-specific reward factor focuses specifically on the tracking performance for each individual task trajectory.
In Deployment: During deployment, the target trajectory is provided either by the pre-trained DP or through human teleoperation using the handheld gripper.
The target trajectory output by the DP is applied to an automated quadruped robot loco-manipulation task. The policy input includes multimodal data such as the arm’s proprioception and corresponding fisheye images at each time step, with the TCP trajectory defining the action space. We train the DP using the ViT-based CLIP visual encoder (ViT Base Patch 16, input 224×224), following the methods in [32] and [33].
Using the handheld gripper for teleoperation is another option that allows humans to directly control the whole-body behaviors of the quadruped robot with arm. As mentioned earlier, the whole-body policy can generalize to track previously unseen trajectories within a reasonable operational range. As a result, we developed a teleoperation scheme to control the quadruped robot, expanding its ability to perform whole-body behaviors.
The Realsense T265 on the handheld device records pose data relative to its origin at startup. We set the robot’s initial TCP point after startup as the origin of its trajectory, ensuring alignment between the initial points of the handheld device and the robot’s TCP. As the operator moves the device, the recorded historical pose data is fed into the policy for tracking, enabling whole-body behaviors under teleoperation.
IV Experiments and Results
IV-A System Setup
During whole-body policy training, we leverage Isaac Gym to collect data in parallel from 4,096 agents. The policy updates every 50 simulation steps, and the entire training–runs on a single NVIDIA RTX 3090 Ti GPU–takes roughly eight hours to complete 10,000 iterations. The policy controller operates at $50\text{\,}\mathrm{H}\mathrm{z}$. We implement asynchronous communication for the DP or teleoperation, whole-body policy, and robot control through multi-thread control, as shown in Figure 3.
Figure 3: System Deployment Architecture: We developed a multi-computer, multi-threaded system for deployment. The host computer runs the DP ($10\text{\,}\mathrm{H}\mathrm{z}$) to generate or capture the handheld gripper’s trajectory and commands ($100\text{\,}\mathrm{H}\mathrm{z}$) using the Realsense T265 and GoPro. The frequency affects task execution speed. This reference trajectory is sent to the lower computer via Ethernet. The Jetson Orin NX, mounted on the robot, handles multi-threaded tasks, including receiving robot status ($500\text{\,}\mathrm{H}\mathrm{z}$), whole-body policy inference ($50\text{\,}\mathrm{H}\mathrm{z}$), and sending control commands ($200\text{\,}\mathrm{H}\mathrm{z}$).
The controller is deployed on a 12-DoF quadruped robot, Uniree Go2, which is equipped with a 6-DoF Airbot Play robotic arm. The arm’s end effector is fitted with a single DoF parallel gripper [11], as shown in Figure 3. The robot total weights $19.9\text{\,}\mathrm{kg}$. Both the arm and the quadruped robot are powered by a battery on Go2. Inference for the whole-body policy runs onboard the Go2’s NVIDIA Jetson Orin NX, while the DP in automated mode is inferred on a PC equipped with an NVIDIA RTX 3090Ti, with the target trajectory sent via Ethernet. In teleoperation mode, a handheld gripper transmits the trajectory and gripper commands over the same link.
During deployment, it is necessary to obtain the target pose and the current TCP pose in the world coordinate system. We install a Livox Mid-360 Lidar on the robot’s body and deploy an odometry [34] to build the world coordinate system, with the initial position of the Lidar serving as the origin. During system operation, the pose of the TCP is calculated through forward kinematics (FK) and transformed into the world coordinate system. The target pose is processed similarly.
IV-B Ablation Studies
We demonstrate the necessity and effectiveness of the proposed method through a series of ablation experiments conducted in simulations. For experiments with the same setup, we executed 20 trials with different seeds to record the data.
Policy network wo TVP: Using the single MLP (sMLP) network-based Actor from [16] as a baseline, we validate the necessity of our TVP policy network for control. In the absence of external perception, quadruped robot movement relies on proprioceptive information to sense the environment. Our training results indicate that, when the component that estimates state using historical proprioceptive information is removed and a single MLP maps the history proprioceptive observation to actions, the task-tracking reward for loco-manipulation tasks rises more slowly during training and converges at a lower level, as shown in Figure 4. Evaluation further shows that this policy causes larger joint torque and position fluctuations, resulting in more pronounced jitter in joint motions, as shown in Figure 5(a) and (b).
Figure 4: The trajectory tracking reward (left) and total reward (right) curve during the training of three policy networks: TVP (Ours), TVP without NAE, and single MLP.
TVP policy network wo NAE: When the future target trajectory is unobservable (teleoperation), we carried out an ablation experiment to assess the performance conferred by the NAE, which predicts forthcoming trajectories from past trajectory data. In this setting, only the historical and current trajectories are encoded, and the resulting latent embedding is fed directly into the low-level network to generate action outputs. Because the policy lacks information about the trajectory it subsequently track, its tracking performance deteriorates: during training, the tracking reward remains lower than that achieved by the policy equipped with the NAE-based prediction mechanism, as illustrated in Figure 4. Figure 5(c) shows the MSE of the future trajectory predicted by the NAE decoder and its ground truth. The reconstruction error is represented by the position and the first two rows of pose matrix for four frames and plotted as a heatmap. The results demonstrate that the reconstruction error is minimal, indicating that the latent embedding of the NAE contains information essential for predicting the future trajectory.
Figure 5: Policy network ablation analysis: (a) Schematic of the ‘unplug charger’ and ‘pushing’ action sequences in simulation. (b) Arm joint torque and position curves for TVP and single MLP networks during the tasks, with sMLP showing more jitter. (c) MSE heatmap of future trajectory reconstruction by TVP’s NAE for both tasks.
Table III shows the policy evaluation for seven tasks. The position and orientation errors of the end-effector’s trajectory tracking were recorded after 800 time steps. Position error is measured using Euclidean distance (cm), while orientation error refers to the minimum rotational angle between the end-effector and the target trajectory’s orientation (radian). The average error over all time steps was computed to assess overall tracking performance. Results show that the TVP, with its future trajectory prediction mechanism, outperforms other policies, while the single MLP-based policy performs the worst due to its limited information extraction and representation ability.
Adaptive trajectory sampling: In this experiment, we set all robots to uniformly sample (Uni-sample) all task trajectories throughout the training process in order to assess the necessity of our adaptive trajectory sampling strategy. We record the reward curves for two typical tasks during training, as shown in Figure 6. The results show that uniform sampling cannot effectively balance task performance. Initially, the reward curves for tasks with different pose variations are close, but over time, the rewards diverge, with tasks requiring more pose variation converging to lower rewards due to insufficient attention (i.e., higher sampling rates). In contrast, adaptive sampling maintains similar rewards across tasks, balancing tracking performance. Table III indicates that tasks with more frequent trajectory changes (e.g., ”open container”) see a significant decline in tracking accuracy.
Figure 6: The trajectory tracking reward for two tasks under adaptive sampling and uniformly sampling.
TABLE III: Simulation ablation study: mean 6D tracking error for each task, and the variance of those mean errors across all tasks. Metrics: Euclidean distance / rotation-matrix angular difference. (cm/radian)
pushing unplug charger pick bear pick cup open container clean table pour coke Variance
TVP (Ours) 1.32/0.08 1.08/0.07 0.87/0.11 0.66/0.05 0.74/0.06 0.94/0.09 1.07/0.15 0.050/0.0012
TVP wo NAE 1.78/0.11 1.49/0.07 1.30/0.13 1.26/0.07 1.43/0.05 1.31/0.10 1.78/0.22 -
sMLP (baseline) 2.23/0.12 2.02/0.14 2.31/0.37 2.24/0.14 2.20/0.13 2.17/0.22 3.13/0.54 -
Uni-sample 1.30/0.07 1.11/0.08 0.96/0.13 0.79/0.06 1.69/0.10 1.18/0.11 1.41/0.21 0.088/0.0026
IV-C Real-world Experiments
We further validate the feasibility and effectiveness of our method through real-world experiments. The experiments include both teleoperation via a handheld gripper and DP-based inference execution. All tasks were accomplished using a unified trained whole-body policy. Our approach demonstrates excellent performance in handling targets across all scenarios, achieving zero-shot transfer from simulation to the real world. Please refer to our videos in the supplemental materials.
Handheld Gripper Teleoperation: Figure 1 I–V showcase several real-world demonstrations of whole-body loco-manipulation using the teleoperation, such as cleaning up paper scraps, picking up trash, storing toys, unplugging chargers, and removing bolts to open doors. Figure 1 III demonstrates a long-range task. It can be observed that when performing ground-level tasks, such as picking up trash, the quadruped robot assists the arm by leaning downward, enabling the end-effector to reach lower positions. Conversely, tasks well above body height require lifting the body. When the target position lies well beyond the current workspace, the quadruped robot exploits its locomotion capability to reposition itself, thereby reaching the distant site. These behaviors demonstrate exceptional whole-body coordination capabilities. We also evaluated the tasks for training and recorded the tracking performance in Table IV. The results show that the real-world deployment achieves tracking performance comparable to that of the simulation.
TABLE IV: Real-world experiment results. (cm/radian/%)
EE pos err EE rot err Success Rate
Teleoperation unplug charger 1.39 0.054 98
open container 1.02 0.047 100
DP unplug charger 1.37 0.039 80
open container 1.24 0.036 85
Inference through DP: As shown in Figure 1 VI–VII, we demonstrate the method’s ability to autonomously execute tasks through a pre-trained DP. Specifically, we input the information from a GoPro camera mounted on the robotic arm’s wrist, along with the arm’s states, into the DP. The DP then autonomously generates the trajectory of the target object. The RL whole-body policy is then used to control the joints for execution. Its performance is mainly limited by the results of DP inference. The experimental results indicate that our method can leverage the DP to plan and autonomously perform multi-task operations, as shown in Table IV.
V Conclusion
We proposed MLM, a reinforcement learning framework that leveraged both real-world and simulation data to address multi-task whole-body loco-manipulation control for quadruped robots. To balance performance while learning multiple task trajectories, we integrate a data library with an adaptive sampling mechanism. To address the scenario where only historical trajectories can be observed during teleoperation, we propose a Trajectory-Velocity Prediction policy network to predict future trajectories, thereby improving policy execution performance. Velocity estimation supports task execution across different spatial ranges. Finally, simulations and real-world experiments on Go2 and Airbot Play robot highlight the good performance of the proposed method. In fact, our approach can be extended to enable quadruped robots to perform more tasks in new environments, fully utilizing their locomotion and manipulation capabilities. This depends on higher-level decision-making and planning. Future integration of vision-language models and navigation could address this limitation.

### v2

@@
 Abstract
-Whole-body loco-manipulation for quadruped robots with arm remains a challenging problem, particularly in achieving multi-task control.
+Whole-body loco-manipulation for quadruped robots with arms remains a challenging problem, particularly in achieving multi-task control.
 To address this, we propose MLM, a reinforcement learning framework driven by both real-world and simulation data.
@@
 By leveraging extensive simulation data and curriculum-based rewards, our controller achieves whole-body behaviors in simulation and zero-shot transfer to real-world deployment.
-Ablation studies in simulation verify the necessity and effectiveness of our approach, while real-world experiments on the Go2 robot with an Airbot robotic arm demonstrate the policy’s good performance in multi-task execution.
-Index Terms:
-Reinforcement learning, legged robot, multi-task loco-manipulation, whole-body control.
+Ablation studies in simulation verify the necessity and effectiveness of our approach, while real-world experiments on a Go2 robot with an Airbot robotic arm demonstrate the policy’s good performance in multi-task execution.
 I Introduction
-Quadruped robot motion control, especially through reinforcement learning (RL), has gained significant attention and development, enhancing their diverse and impressive capabilities [1, 2, 3, 4, 5, 6].
+Quadruped robot motion control, especially through reinforcement learning (RL), has gained significant attention and development, enhancing their diverse and impressive capabilities [lee2020learning, wjzamp, kumar2021rma, ji2022concurrent, Li2024Learning, Peng2025Learning].
 A six-degree-of-freedom (DoF) robotic arm can offer more possibilities for enabling quadruped robots to perform manipulations.
-By integrating the locomotion of a quadruped robot with the manipulation capabilities of a robotic arm, a broader range of tasks can be accomplished compared to using a quadruped robot alone [7, 8].
-However, achieving multiple loco-manipulation tasks with a single policy presents a challenge for quadruped robots.
-Employing a single policy across multiple tasks can reduce control complexity by eliminating the need for task-specific policies.
-However, balancing multiple tasks remains a problem that has not been adequately addressed by existing methods.
-Consequently, an adaptive balancing mechanism is essential.
-The method for achieving adaptive balance requires further research.
-In simulation-driven RL paradigms, the virtual assets and rewards specific to different tasks often vary, and careful, personalized design typically leads to a significant workload and lacks generalizability.
-Furthermore, the gap of visual information required for manipulation is often difficult to address due to significant differences in texture, color, lighting, and other factors between simulation and the real world [9].
+By integrating the locomotion of a quadruped robot with the manipulation capabilities of a robotic arm, a broader range of tasks can be accomplished [Sleiman2023Versatile, Yokoyama2024ASC].
+Figure 1: The robot showcases its ability to perform multi-task whole-body loco-manipulation in the real world.
+Employing a single policy can reduce control complexity by eliminating the need for task-specific policies.
+However, achieving multiple loco-manipulation tasks with a single policy presents a challenge for quadruped robots, and balancing multiple tasks remains unresolved in existing methods, making an adaptive mechanism essential and warranting further research.
+Moreover, in simulation-driven paradigms, task-specific assets and rewards demand heavy manual design with poor generalizability.
+The visual gap for manipulation also remains challenging due to differences in texture, color, lighting, and other factors between simulation and reality [Mohammad2024SplatSim].
 A feasible solution is to distill multiple tasks into multiple trajectories learning.
-This avoids incorporating observational data of the task objectives (such as visual information) into the control, instead replacing it with trajectories.
-Figure 1: The robot showcases its ability to perform multi-task whole-body loco-manipulation in the real world.
-The acquisition of trajectories can be achieved through various methods, including UMI-like approaches [10, 11], teleoperation [12, 13], exoskeleton-based systems [14].
-It is important to note that we do not collect whole-body trajectory data for quadruped loco-manipulation, which distinguishes our approach from imitation learning.
-This is due to the challenges of species selection for data collection and the inability to obtain such data from real robots.
+This avoids incorporating visual data of the task objectives into the control.
+Building on this idea, real-world multi-task trajectories can be introduced into simulation, while simulation can generate large-scale interaction data.
+By abstracting tasks into trajectories and defining a unified reward, quadruped robots can be guided to develop coordinated whole-body behaviors under a single RL policy.
+Real-world multi-task data provide spatial trajectories consistent with physical dynamics and distributed across the robot’s work space, which are essential for learning whole-body behaviors and generalizing to unseen trajectories [He2025ASAP].
+In contrast, random spatial curves sampled in simulation lack realistic dynamics such as velocity and acceleration.
+The acquisition of trajectories can be achieved through various methods, including UMI-like approaches [chi2024umi, Zhaxizhuoma2025FastUMI], teleoperation [Tony2023Learning, Anthony2023RT1], exoskeleton-based systems [Fang2024AirExo].
+It is important to note that we do not collect the whole-body trajectory data, which distinguishes our approach from imitation learning.
+This is due to the challenges of species selection for data collection and the inability to obtain such data from real quadruped robot with an arm.
 Thus, we use the UMI-like approach to obtain only the end-effector 6D manipulation trajectories, a lightweight solution independent of the robots.
-Integrating real-world multi-task manipulation trajectories into simulation and leveraging the simulation to efficiently generate training data that interacts with the environment is a feasible approach.
-Furthermore, by abstracting multiple tasks into multiple trajectories and designing a unified task reward, it can guide the quadruped robot in generating coordinated whole-body behaviors under a single policy to accomplish the tasks.
-These multi-task data from the real world provide spatial trajectories that align with the dynamics of the physical environment and are distributed across the robot’s manipulation space, which is essential for learning whole-body behaviors and generalizing to unseen trajectories [15].
-This contrasts with sampling random spatial curves in simulation, which lack real-world dynamics like velocity and acceleration.
-It is important to note that in deployment schemes that only have access to historical trajectories, such as teleoperation, future trajectory information cannot be used as observation.
-However, future information is crucial for the stability of whole-body execution tasks, as it indicates the direction of the next step in the action [16].
+In deployment schemes that only have access to historical trajectories, such as teleoperation, future trajectory information cannot be used as observation.
+However, future information is crucial for the stability of whole-body execution tasks, as it indicates the direction of the next step in the action [Huy2024UMIonLegs].
 Thus, a trajectory prediction mechanism is necessary.
-To address the aforementioned challenges, we introduce MLM, a learning-based multi-task loco-manipulation method.
-We integrate Fast-UMI [11] to provide manipulation trajectories.
-In the simulation, we train the control policy and transfer it to the real world.
+To address the aforementioned challenges, we incorporate Fast-UMI [Zhaxizhuoma2025FastUMI] to provide real-world manipulation trajectories and introduce MLM, a learning-based multi-task loco-manipulation framework.
+The control policy is trained in simulation and subsequently transferred to the real world.
 Our contributions are summarized as follows:
-  • We propose a Trajectory-Velocity Prediction-based framework for training a multi-task loco-manipulation controller, driven by both real-world and simulation data, with trajectory interface deployment enabled through teleoperation and a diffusion policy.
-  • We introduce a multi-task real-world trajectory data library, where adaptive and curriculum-based sampling enables the robot to balance performance across tasks and learn multi-task manipualtion with a single policy network.
-  • We deploy a real quadruped robot with arm and validate the excellent performance of our method in whole-body loco-manipulation control across multiple tasks.
+  • We propose a Trajectory-Velocity Prediction-based RL framework for training a multi-task loco-manipulation controller, driven by both real-world and simulation data, with trajectory interface deployment enabled through teleoperation and a diffusion policy.
+  • We introduce a multi-task real-world trajectory data library, where adaptive and curriculum-based sampling enables the robot to balance performance across tasks and learn multi-task manipulation with a single policy network.
+  • We deploy a real quadruped robot with an arm and validate the excellent performance of our method in whole-body loco-manipulation control across multiple tasks.
 II Related Work
 II-A RL-based Legged Locomotion
-The success of RL-based motion controllers for legged robots, compared to model-based approaches, stems from the following key advantages.
-RL trains control policies using large amounts of data collected from simulated robot-environment interactions, capturing the dynamics of these interactions while leveraging far more data than real-world training allows [1, 3, 17].
-Furthermore, reward-driven policy networks provide an end-to-end mapping from states to joint actions, eliminating the need for costly optimization during inference [18, 4].
-Building on the parallel dynamic simulator Isaac Gym, a quadruped robot was trained to walk within minutes [19].
-An actuator network was employed to model the complex actuator dynamics of the ANYmal robot, thereby enabling sim-to-real transfer [20].
-Extending this approach, privileged learning was introduced to train the ANYmal robot, successfully achieving a blind motion controller for complex terrains [1].
-Additionally, a method named RMA was proposed to develop robust quadruped locomotion controllers [3].
-To facilitate gait learning and improve adaptation to complex terrains, gait priors were introduced [21, 2, 22], allowing robots to leverage motion data from flat terrains and rapidly converge to the desired gait pattern.
-While these efforts have significantly advanced the locomotion capabilities of legged robots [23, 24, 25, 26], the development of their loco-manipulation abilities still requires further exploration.
+The success of RL-based motion controllers for legged robots, compared to model-based methods, lies in two main advantages.
+First, RL trains control policies using large-scale simulated robot–environment interactions, capturing robot dynamics while leveraging far more data than real-world training allows [lee2020learning, kumar2021rma, miki2022learning].
+Second, reward-driven policies enable end-to-end mapping from states to joint action, eliminating the need for costly optimization during inference [margolis2022rapid, ji2022concurrent].
+Built on the parallel simulator Isaac Gym, quadruped robots can learn to walk within minutes [rudin2022learning].
+To enable sim-to-real transfer, an network was used to capture the complex dynamics of the ANYmal robot’s motors [hwangbo2019learning].
+This approach was later extended with privileged learning, yielding a blind locomotion controller for challenging terrains [lee2020learning].
+Further advances include RMA, which enhances robustness in quadruped locomotion [kumar2021rma].
+To accelerate gait learning and adapt to complex terrains, researchers introduced gait priors [escontrela2022adversarial, wjzamp, Liu2024Skill], leveraging flat-terrain motion data for faster convergence.
+These efforts have greatly advanced legged locomotion [Fan2024Rethinking, David2024ANYmalparkour, Miki2024Learning3D, Tairan2024Agile], yet the development of loco-manipulation remains largely unexplored.
 II-B Legged Loco-Manipulation
-To endow the manipulation capabilities of legged robots, researchers have integrated the mobility of legged platforms with the manipulation abilities of robotic arms.
-For instance, in [27], footstep planning was addressed, and a hierarchical optimization-based controller was developed to enable a quadruped robot with a six-DoF robotic arm to perform manipulation.
-The study in [7] proposed a approach that searches for whole-body trajectories and determines contact schedules to solve mobile manipulation tasks within predefined environments.
+To enhance legged robots with manipulation skills, researchers have combined their mobility with the manipulation capabilities of robotic arms.
+For instance, in [Bellicoso2019ALMA], footstep planning was addressed, and a hierarchical optimization-based controller was developed to enable a quadruped robot with a six-DoF robotic arm to perform manipulation.
+The study in [Sleiman2023Versatile] proposed a approach that searches for whole-body trajectories and determines contact schedules to solve mobile manipulation tasks within predefined environments.
 Nonetheless, these optimization-based approaches often require task-specific design, presenting significant challenges in integrating the execution of multiple tasks within a single control policy.
 Additionally, they incur substantial computational overhead due to the optimization processes involved.
-In recent years, several studies have focused on employing RL to achieve end-to-end control of legged robots with robotic arms and exploring the execution of policies across multiple tasks.
-In [28], the arm’s motion is modeled as predictable external torques that the RL-trained locomotion policy uses to compensate for the manipulator’s influence.
-Nevertheless, this non-whole-body approach limits the ability of the arm to utilize the mobility and posture adjustment capabilities of the quadruped robot to expand its manipulation range.
-To achieve whole-body control, an RL-based Regularized Online Adaptation method was proposed, successfully achieving sim-to-real transfer for quadruped loco-manipulation [29].
-Despite this, the manipulation space defined under the arm base are susceptible to the influence of body shaking and are not stable enough.
-We utilize world space to address this issue, as the operational objectives are independent of body movement.
-In [30], a nonlinear reward fusion module was proposed to achieve whole-body behaviors for a wheeled quadrupedal manipulator.
+In recent years, several studies have focused on employing RL to achieve end-to-end control of legged robots with robotic arms.
+In [Ma2022Combining], the arm’s motion is modeled as predictable external torques, which the RL-trained locomotion policy compensates for.
+However, this non-whole-body approach limits the arm from leveraging the quadruped’s mobility and posture adjustments to extend its manipulation range.
+To enable whole-body control, an RL-based regularized online adaptation method was proposed, successfully achieving sim-to-real transfer for quadruped loco-manipulation [Fu2023DeepWholeBody].
+Despite this, the arm-base-defined manipulation space is vulnerable to body shaking and thus unstable.
+To address this, we adopt a world frame, where operational objectives remain independent of body motion.
+In [Jiang2025Learning], a nonlinear reward fusion module was proposed to achieve whole-body behaviors for a wheeled quadrupedal manipulator.
 However, it did not integrate automatic execution of the task.
-In [31], an RL-based approach that integrates visual information for autonomous loco-manipulation execution was proposed.
+In [Liu2024VisualWholeBody], an RL-based approach that integrates visual information for autonomous loco-manipulation execution was proposed.
 Yet, the inverse kinematics (IK) strategy is also susceptible to locomotion interference due to the arm base being affected by the body’s movement.
-In [16], a vision-driven, trajectory-based loco-manipulation in the task space was realized for a quadruped robot.
+In [Huy2024UMIonLegs], a vision-driven, trajectory-based loco-manipulation in the task space was realized for a quadruped robot.
 However, it did not consider the integration of multiple tasks and the balance of performance across tasks within a single policy.
@@
 The pipeline is illustrated in Figure 2.
-Figure 2: Overview of the training and inference pipeline.
-We use reinforcement learning to train a whole-body control policy.
-A trajectory library, together with an adaptive sampling curriculum, supplies a multiple set of real-world manipulation trajectories for the policy to track.
-A normalized autoencoder encodes historical target trajectories, implicitly conveying information about upcoming targets; alternatively, the policy can receive future trajectories directly, enabling two deployment modes: teleoperation and DP generation.
-We also train an estimator to estimate base linear velocity $\hat{v}_{t}^{p}$ using supervised learning.
+Figure 2: Overview of the pipeline.
+We use RL to train a whole-body control policy.
+A trajectory library, together with an adaptive sampling curriculum, supplies multiple real-world manipulation trajectories for the policy to track.
+An NAE encodes historical trajectories to estimate future targets, while the policy can also take future trajectories directly, supporting two deployment modes: teleoperation and automatic DP.
+In addition, we train a supervised estimator to predict the base linear velocity $\hat{v}_{t}^{p}$.
 III-A Whole-body Loco-Manipulation Policy
-We eliminate the requirement for object and environment perception in simulation, which speeds up training and avoids the large sim-to-real gap caused by visual perception.
-Since information is partially observable, we model the loco-manipulation task as a partially observable Markov decision process (POMDP) [17, 22].
-We employ RL to train a leg-arm whole-body policy $\pi_{\theta}$, aiming to find optimal policy parameters $\theta$ that maximize the discounted expected return $\mathbb{E}_{\mathbf{\tau}(\pi_{\theta})}[\sum_{t=0}^{T}{\gamma^{t}r_{t}}]$, where $r_{t}$ is the reward at time $t$, and $\gamma$ is the discount factor ($\gamma=0.99$).
-We introduce an asymmetric Actor-Critic (AAC) framework to train the loco-manipulation policy.
-This approach is designed to address the POMDP problem while also avoiding the complexity of the two-stage training process [4].
-State Space: The full state for the policy includes the quadruped robot’s proprioception $\boldsymbol{o}_{t}^{q}\in{\mathbb{R}^{42}}$, the six-DoF robotic arm’s proprioception $\boldsymbol{o}_{t}^{a}\in{\mathbb{R}^{18}}$, the trajectory pose sequence $o_{t-3:t+4}^{tcp}\in{\mathbb{R}^{72}}$, the privileged state $\boldsymbol{s}_{t}^{p}\in{\mathbb{R}^{21}}$, and the terrain elevation data $\boldsymbol{s}_{t}^{e}\in{\mathbb{R}^{187}}$.
-Specifically, $\boldsymbol{o}_{t}^{q}$ includes the quadruped’s body angular velocity $\boldsymbol{\omega}_{t}^{q}\in{\mathbb{R}^{3}}$, the gravity vector $\boldsymbol{g}_{t}^{q}\in{\mathbb{R}^{3}}$ representing its orientation, leg joint positions $\boldsymbol{\theta}_{t}^{q}\in{\mathbb{R}^{12}}$ and velocities $\dot{\boldsymbol{\theta}}_{t}^{q}\in{\mathbb{R}^{12}}$, and the previous time step’s leg joint actions $\boldsymbol{a}_{t-1}^{q}\in{\mathbb{R}^{12}}$.
-The arm proprioception $\boldsymbol{o}_{t}^{a}$ contains the arm joint positions $\boldsymbol{\theta}_{t}^{a}\in{\mathbb{R}^{6}}$, velocities $\dot{\boldsymbol{\theta}}_{t}^{a}\in{\mathbb{R}^{6}}$, and last arm joint actions $\boldsymbol{a}_{t-1}^{a}\in{\mathbb{R}^{6}}$.
+We eliminate the requirement for object and environment perception in simulation, which accelerates training and avoids the large sim-to-real visual gap.
+Since information is partially observable, the loco-manipulation task is modeled as a partially observable Markov decision process (POMDP) [miki2022learning].
+RL is employed to train a whole-body policy $\pi_{\theta}$, aiming to find optimal parameters $\theta$ that maximize the discounted expected return $\mathbb{E}_{\mathbf{\tau}(\pi_{\theta})}[\sum_{t=0}^{T}{\gamma^{t}r_{t}}]$, where $r_{t}$ is the reward at time $t$, and $\gamma$ is the discount factor ($\gamma=0.99$).
+We introduce an asymmetric actor-critic framework to train the policy [ji2022concurrent].
+This approach is designed to address the POMDP problem while also avoiding the complexity of the two-stage training.
+State space: The input of the actor (policy network) includes a five-frame history sequence of the quadruped robot’s proprioception $\boldsymbol{o}_{t}^{q}\in{\mathbb{R}^{30}}$, the six-DoF robotic arm’s proprioception $\boldsymbol{o}_{t}^{a}\in{\mathbb{R}^{12}}$, the last action $\boldsymbol{a}_{t-1}\in{\mathbb{R}^{18}}$, and the end-effector pose trajectory $o_{t-3:t+4}^{tcp}\in{\mathbb{R}^{72}}$.
+Besides the actor’s input, the critic has access to the privileged state $\boldsymbol{s}_{t}^{p}\in{\mathbb{R}^{21}}$ and the terrain elevation data $\boldsymbol{s}_{t}^{e}\in{\mathbb{R}^{187}}$, which are only available during training and not used by the actor.
+A normalized autoencoder (NAE) only receives the historical end-effector pose sequence to predict future trajectory, while the memory encoder only receives the historical robot and arm proprioception.
+Further details are provided in Section III-B.
+Specifically, $\boldsymbol{o}_{t}^{q}$ includes the quadruped’s body angular velocity $\boldsymbol{\omega}_{t}^{q}\in{\mathbb{R}^{3}}$, the gravity vector $\boldsymbol{g}_{t}^{q}\in{\mathbb{R}^{3}}$ representing its orientation, leg joint positions $\boldsymbol{\theta}_{t}^{q}\in{\mathbb{R}^{12}}$ and velocities $\dot{\boldsymbol{\theta}}_{t}^{q}\in{\mathbb{R}^{12}}$.
+The arm proprioception $\boldsymbol{o}_{t}^{a}$ contains the arm joint positions $\boldsymbol{\theta}_{t}^{a}\in{\mathbb{R}^{6}}$ and velocities $\dot{\boldsymbol{\theta}}_{t}^{a}\in{\mathbb{R}^{6}}$.
+The last action $\boldsymbol{a}_{t-1}$ includes the previous time step’s leg joint action $\boldsymbol{a}_{t-1}^{q}\in{\mathbb{R}^{12}}$ and arm joint action $\boldsymbol{a}_{t-1}^{a}\in{\mathbb{R}^{6}}$.
 The trajectory pose sequence $o_{t-3:t+4}^{tcp}$ comprises Tool Center Points (TCP) over past and subsequent steps relative to time $t$, each including a 3D position and the first six elements of the rotation matrix.
-The privileged state $\boldsymbol{s}_{t}^{p}$, covering both robot and environment, includes the quadruped’s body linear velocity $\boldsymbol{v}_{t}^{q}\in{\mathbb{R}^{3}}$, contact forces $\boldsymbol{f}^{c}_{t}\in{\mathbb{R}^{12}}$, external force $\boldsymbol{f}^{e}_{t}\in{\mathbb{R}^{3}}$, and its position $\boldsymbol{p}^{e}_{t}\in{\mathbb{R}^{3}}$ on the robot.
-Finally, $\boldsymbol{s}_{t}^{e}\in{\mathbb{R}^{187}}$ represents the heights of 187 terrain points evenly distributed around the robot’s coordinate frame.
-Action Space: The whole-body policy outputs 12 joint position offsets for the legs, $\boldsymbol{a}_{t}^{q}\in{\mathbb{R}^{12}}$, and 6 joint position offsets for the arm, $\boldsymbol{a}_{t}^{a}\in{\mathbb{R}^{6}}$, both relative to their default positions.
-These offsets are added to the default positions and sent to the motors as target position commands for proportional-derivative (PD) control [30].
-Reward: Maintaining coordinated loco-manipulation of quadruped robots is challenging.
-Building on [22], we use adversarial motion prior (AMP) to encourage quadruped gaits learning via style rewards.
-During policy training, an adversarial discriminator network is trained concurrently to provide style rewards, defined as $r^{s}=\max[0,1-0.25\times\left(d^{\rm score}-1\right)^{2}]$.
+The privileged state $\boldsymbol{s}_{t}^{p}$, covering both robot and environment, includes the quadruped’s body linear velocity $\boldsymbol{v}_{t}^{q}\in{\mathbb{R}^{3}}$, contact force $\boldsymbol{f}^{c}_{t}\in{\mathbb{R}^{12}}$, external force $\boldsymbol{f}^{e}_{t}\in{\mathbb{R}^{3}}$, and its position $\boldsymbol{p}^{e}_{t}\in{\mathbb{R}^{3}}$ on the robot.
+Finally, $\boldsymbol{s}_{t}^{e}$ represents the heights of 187 terrain points evenly distributed around the robot frame.
+Action space: The whole-body policy outputs 12 joint position offsets for the legs, $\boldsymbol{a}_{t}^{q}\in{\mathbb{R}^{12}}$, and 6 joint position offsets for the arm, $\boldsymbol{a}_{t}^{a}\in{\mathbb{R}^{6}}$.
+These offsets are added to the default positions and sent to the motors as target position commands for proportional-derivative (PD) control [Jiang2025Learning].
+Reward: Maintaining coordinated whole-body behaviors is challenging.
+Building on [Liu2024Skill], we use adversarial motion prior (AMP) to encourage quadruped gaits learning via a style reward.
+During policy training, an adversarial discriminator is trained concurrently to provide the style reward, defined as $r^{s}=\max[0,1-0.25\times\left(d^{\rm score}-1\right)^{2}]$.
 The primary reward of the policy, $r^{g}$, is the tracking of the robotic arm’s TCP position and orientation:
@@
 $\oplus$’ represents the concatenation operation.
-We have established a curriculum to update different tasks reward factor $\sigma_{n}^{p}$ and $\sigma_{n}^{o}$, more details are described in III-C.
-The constraint reward, $r^{l}$, is used to regulate the robot’s reasonable actions, energy limitations, and hardware constraints.
+We have established a curriculum to update different tasks reward factors $\sigma_{n}^{p}$ and $\sigma_{n}^{o}$, more details are described in III-C.
+The constraint reward, $r^{l}$, is used to regulate the robot’s reasonable action, energy limitations, and hardware constraints.
 All rewards are shown in Table I.
@@
 $r^{s}$ $\max[0,1-0.25\times(d^{\rm score}-1)^{2}]$ 0.5
-Terrain Curriculum: In addition to the enhanced ability to adjust its posture with more degrees of freedom, the leg configuration of quadruped robots enables flexible adaptation to complex terrains.
-To enable a quadruped robot to perform loco-manipulation tasks in complex terrains, we introduce terrain curriculum.
-This approach allows the robot to progressively learn whole-body behaviors across several types of terrain, starting from easier to more difficult conditions.
-These terrains include randomly undulating surfaces, slopes, discrete obstacles, stairs.
-Unlike pure locomotion tasks, we have removed the terrain traversal update mechanism, and instead, we use reward threshold updates.
-Specifically, we monitor the TCP tracking reward within an episode, and when it reaches the threshold, the robot progresses to the next level of the curriculum in the subsequent episode; otherwise, it regresses to the previous level.
-We have found that this mechanism effectively ensures the robot maintains a good tracking performance while preserving the stability of the base.
-Domain Randomization: To reduce the sim-to-real gap, we randomize robot hardware, control, and environment parameters within specified ranges during training to ensure the simulation covers or approximates real-world conditions.
+Terrain curriculum: In addition to offering greater freedom of posture adjustment, the leg design of quadruped robots enables adaptation to terrains.
+To support loco-manipulation in such environments, we introduce a terrain curriculum [Liu2024Skill], which gradually trains the robot to master whole-body behaviors, progressing from simple to challenging terrains.
+These terrains include randomly undulations, slopes, discrete obstacles, and stairs.
+Unlike locomotion tasks, we remove the terrain traversal update mechanism, and instead apply reward threshold updates.
+Specifically, we monitor the tracking reward within an episode, and when it reaches the threshold, the robot progresses to the next curriculum level in the subsequent episode; otherwise, it regresses to the previous one.
+This mechanism effectively maintains good tracking performance while ensuring base stability.
+Domain randomization: To reduce the sim-to-real gap, we randomize robot hardware, control, and environment parameters within specified ranges during training [wjzamp].
 As shown in Table II.
@@
 III-B Network
-The training framework consists of a policy network and a Critic network. The policy network acts as the Actor and controller, receiving only observable states as input.
-Trajectory-Velocity Prediction Policy Network: We propose a Trajectory-Velocity Prediction Policy Network (TVP), which predicts unobservable target trajectories and body velocity, integrates additional proprioception, and outputs joint actions.
-A trajectory encoder $E_{\theta_{1}}$ extracts features from the historical pose frames.
+Trajectory-velocity prediction policy network: We propose a Trajectory-Velocity Prediction policy network (TVP) consisting of an NAE, a memory encoder, an estimator and a low-level network.
+The historical trajectories are processed by the NAE to first obtain the NAE embeddings and then predict future poses.
+The historical proprioceptive state sequence is fed into the memory encoder (ME) to produce the ME embeddings.
+The ME embeddings are used as inputs to the estimator, which predicts the body’s linear velocity.
+The predicted future poses, the NAE embeddings, the predicted linear velocity, the ME embeddings, the current proprioceptive state and the last action are fed into a low-level network to generate the joint action.
+A trajectory encoder $E_{\theta_{1}}$ extracts features from the historical poses.
 Then, these features are fed into a decoder $D_{\theta_{2}}$ to predict the subsequent trajectory sequence, as shown in (2).
-This prediction is essential, particularly in scenarios such as teleoperation, where only historical trajectory records are available.
+This prediction is essential, particularly in teleoperation, where only historical trajectory records are available.
 Alternatively, when deploying the strategy that utilize DP, the future trajectory can be provided directly, and the TVP will be skipped.
@@
 We introduce a normalized autoencoder (NAE) to formulate the structure and loss function for the trajectory predictor.
-The reconstruction loss measures how well the NAE reconstructs the future trajectory from the latent embedding $\hat{z}_{t}^{tcp}$:
+The reconstruction loss measures how well the NAE reconstructs the future trajectory from the embeddings $\hat{z}_{t}^{tcp}$:
 $\mathcal{L}_{NAE}$ $=\mathcal{L}_{rec}=\mathrm{MSE}\left(\hat{o}_{t+1:t+4}^{tcp},o_{t+1:t+4}^{tcp}\right)$ (3)
 Specifically, the MSE is used to minimize the discrepancy between the predicted future trajectory and the ground truth.
-Importantly, we apply the tanhtanh activation function to the encoded latent embedding for normalization, constraining them within a limited range.
+Importantly, we apply the tanhtanh activation function to the encoded embeddings for normalization, constraining them within a limited range.
 This helps prevent outliers that lie outside the expected distribution, thus aiding in stabilizing the training process.
-Additionally, we update the trajectory predictor, constructed by NAE, concurrently with other networks.
-On one hand, this facilitates accurate prediction of future trajectories; on the other hand, it generates useful latent embedding for use by the low-level network, thereby promoting the learning of the desired actions.
+Additionally, we update the trajectory predictor, concurrently with other networks.
+This both facilitates accurate prediction of future trajectories and provides useful embeddings for the low-level network, thereby promoting the learning of the desired action.
 The memory encoder encodes the last five frames of proprioceptive observations. And an estimator is trained through supervised learning to estimate the body’s linear velocity:
@@
 The estimated linear velocity of the body effectively captures the distinction between these two types of tasks, enabling more accurate execution of these different loco-manipulation tasks.
-The encoded embedding, along with other proprioceptive data and previous actions, is fed into a low level network, which then outputs the mean vector $\boldsymbol{\mu}_{t}^{b}\in{\mathbb{R}^{18}}$ of a Gaussian distribution $\boldsymbol{a}_{t}^{b}\sim\mathcal{N}\left(\boldsymbol{\mu}_{t}^{b},\boldsymbol{\sigma}^{b}\right)$, where $\boldsymbol{\sigma}^{b}\in{\mathbb{R}^{18}}$ represents the variance of the action, optimized by proximal policy optimization (PPO).
+The encoded embeddings and predicted values, together with current proprioceptive state and the last action, are fed into a low level network, which then outputs the mean vector $\boldsymbol{\mu}_{t}^{b}\in{\mathbb{R}^{18}}$ of a Gaussian distribution $\boldsymbol{a}_{t}^{b}\sim\mathcal{N}\left(\boldsymbol{\mu}_{t}^{b},\boldsymbol{\sigma}^{b}\right)$, where $\boldsymbol{\sigma}^{b}\in{\mathbb{R}^{18}}$ represents the variance of the action, optimized by proximal policy optimization (PPO).
 Critic network: The critic network comprises a trajectory encoder $E_{\psi_{1}}$ and a height encoder $E_{\psi_{2}}$, which perform dimensionality reduction and representation.
 $\psi_{1}$ and $\psi_{2}$ represent the critic network parameters.
-These encoded features, combined with proprioceptive observations, previous actions and privilege information, are used to compute the state value function $V_{t}$, assessing the expected reward for the current state:
+These encoded features, combined with proprioceptive observations, last action and privilege information are used to compute the state value function $V_{t}$, assessing the expected reward for the current state:
 $V_{t}=E_{\psi_{3}}\left(E_{\psi_{1}}(o_{t-3:t+4}^{tcp})\oplus E_{\psi_{2}}(s_{t}^{e})\oplus o_{t}\oplus a_{t-1}\oplus s_{t}^{p}\right)$ (5)
 III-C Trajectory Generation and Sampling
-Trajectory generation encompasses both the trajectories learned during training and those tracked during deployment.
-The following details the generation and sampling strategies for these trajectories.
-Trajectory Library in Training: We incorporate a trajectory library in whole-body policy training, which contains real-world TCP pose data from multiple manipulation tasks.
+Trajectory generation encompasses both training and deployment, with the subsequent sections detailing the corresponding generation and sampling strategies.
+Trajectory library in training: We introduce a trajectory library containing real-world TCP poses from multiple manipulation tasks.
 During training, the initial points of these trajectories are set as the end-effector poses following episodes reset, preventing low-quality samples in the early stages.
-We select trajectories from several typical manipulation tasks for learning [11].
-Each task includes 200 trajectories, such as ‘clean table’, ‘pour coke’, ‘open container’, ‘pick bear’, ‘pick cup’, and ‘unplug charger’.
+We select trajectories from several typical tasks for learning, such as ‘clean table’, ‘pour coke’, ‘open container’, ‘pick bear’, ‘pick cup’, and ‘unplug charger’ [Zhaxizhuoma2025FastUMI].
+Each task includes 200 trajectories.
 More trajectories can be added to expand multi-task learning.
-We also incorporate ‘pushing’ data from [16].
-These tasks cover repetitive motions, pick-and-place actions, rotational opening and closing, and long-distance movements, making them representative of common real-world operational tasks.
-The data are stored as separate pickle files for training.
-Furthermore, we find that with the guidance of the TCP trajectory tracking reward $r^{g}$, the robot can generalize to tracking other unseen trajectories to some extent.
+We also incorporate ‘pushing’ from [Huy2024UMIonLegs].
+These tasks cover repetitive motions, pick-and-place, rotational opening and closing, and long-distance movements, making them representative of common real-world manipulation tasks.
+The data is stored as separate pickle files.
+Furthermore, guided by the tracking reward, the robot demonstrates a certain ability to generalize to previously unseen trajectories.
 Trajectory sampling update: The complexity of task trajectories varies: some involve more changes in both position and orientation, while others involve fewer changes.
@@
 $P_{n}=\frac{P^{\prime}_{n}}{{\textstyle\sum_{i=1}^{N}}P^{\prime}_{i}},i,n=1,.,N$
-where $\bar{r}_{n}^{g}$ denotes the average tracking reward for each task trajectory. To prevent forgetting, this mechanism retains a minimum probability for each trajectory.
+where $\bar{r}_{n}^{g}$ denotes the average tracking reward for each task trajectory.
+$\lambda_{n}$ is the reward performance threshold for task $n$.
+By tuning the values of $\lambda_{n}$, we can adjust the sampling probability curves of the reference trajectories of different tasks, so that we can achieve similar reward levels for different tasks.
+$\lambda_{n}$ is determined by the user according to reward curves of all tasks after policy training.
+To prevent forgetting, this mechanism retains a minimum probability for each trajectory.
 Meanwhile, each task trajectory adopts its own reward coefficient factor.
-Specifically, after repeated training and adjustments, a predefined set of average pose error thresholds and task reward factors is established as $\left\{\left(T_{k,n},\sigma_{k,n}\right)\right\}$.
-The average pose error for each task is calculated as follows, including Euclidean distance in displacement space and angular error in rotation space.
+Specifically, after repeated adjustments, a predefined set of average pose error thresholds and task reward factors is established as $\left\{\left(T_{k,n},\sigma_{k,n}\right)\right\}$.
+The average pose error for each task is calculated as follows, including Euclidean distance and angular error.
 $e_{n}^{p}=\|\boldsymbol{p}_{n,{\rm tcp}}^{\rm target}-\boldsymbol{p}_{n,{\rm tcp}}\|_{2}$ (7)
@@
 […]
-The factor is adaptively updated based on changes in the reward values for each task trajectory.
-As the training progresses, this factor increases the tracking requirements for each task.
-The task-specific reward factor focuses specifically on the tracking performance for each individual task trajectory.
-In Deployment: During deployment, the target trajectory is provided either by the pre-trained DP or through human teleoperation using the handheld gripper.
-The target trajectory output by the DP is applied to an automated quadruped robot loco-manipulation task.
-The policy input includes multimodal data such as the arm’s proprioception and corresponding fisheye images at each time step, with the TCP trajectory defining the action space.
-We train the DP using the ViT-based CLIP visual encoder (ViT Base Patch 16, input 224×224), following the methods in [32] and [33].
-Using the handheld gripper for teleoperation is another option that allows humans to directly control the whole-body behaviors of the quadruped robot with arm.
-As mentioned earlier, the whole-body policy can generalize to track previously unseen trajectories within a reasonable operational range.
-As a result, we developed a teleoperation scheme to control the quadruped robot, expanding its ability to perform whole-body behaviors.
-The Realsense T265 on the handheld device records pose data relative to its origin at startup.
-We set the robot’s initial TCP point after startup as the origin of its trajectory, ensuring alignment between the initial points of the handheld device and the robot’s TCP.
-As the operator moves the device, the recorded historical pose data is fed into the policy for tracking, enabling whole-body behaviors under teleoperation.
+The factor is adaptively updated based on changes in the reward values for each task trajectory. As the training progresses, this factor increases the tracking requirements for each task.
+In deployment: During deployment, the target trajectory can be provided either by teleoperation or by pre-trained DP.
+Using the handheld gripper for teleoperation allows humans to directly control the quadruped robot with an arm, and expands the robot’s ability to perform whole-body behaviors.
+As mentioned earlier, the policy can generalize to track previously unseen trajectories within a reasonable range.
+The handheld device uses the RealSense T265 to record poses.
+At startup, we set the arm’s TCP as the trajectory origin, aligning it with the T265’s initial point.
+As the operator moves, the recorded poses are fed into the policy for tracking.
+Meanwhile, GoPro-based marker distance estimation directly triggers gripper commands.
+The target trajectory output by the DP is applied to an automated loco-manipulation task.
+In our case, the DP is a pre-trained module using the training approach in [Chi2023Diffusion, Zhaxizhuoma2025FastUMI].
+In the DP training, a ViT-based CLIP visual encoder [Radford2021Learning] is employed with ViT Base/16 architecture and an image resolution of 224×224.
+The policy input includes the arm’s proprioception and fisheye images at each time step, with the target trajectory defining the output.
 IV Experiments and Results
 IV-A System Setup
-During whole-body policy training, we leverage Isaac Gym to collect data in parallel from 4,096 agents.
-The policy updates every 50 simulation steps, and the entire training–runs on a single NVIDIA RTX 3090 Ti GPU–takes roughly eight hours to complete 10,000 iterations.
-The policy controller operates at $50\text{\,}\mathrm{H}\mathrm{z}$.
-We implement asynchronous communication for the DP or teleoperation, whole-body policy, and robot control through multi-thread control, as shown in Figure 3.
-Figure 3: System Deployment Architecture: We developed a multi-computer, multi-threaded system for deployment.
-The host computer runs the DP ($10\text{\,}\mathrm{H}\mathrm{z}$) to generate or capture the handheld gripper’s trajectory and commands ($100\text{\,}\mathrm{H}\mathrm{z}$) using the Realsense T265 and GoPro.
-The frequency affects task execution speed.
-This reference trajectory is sent to the lower computer via Ethernet.
-The Jetson Orin NX, mounted on the robot, handles multi-threaded tasks, including receiving robot status ($500\text{\,}\mathrm{H}\mathrm{z}$), whole-body policy inference ($50\text{\,}\mathrm{H}\mathrm{z}$), and sending control commands ($200\text{\,}\mathrm{H}\mathrm{z}$).
-The controller is deployed on a 12-DoF quadruped robot, Uniree Go2, which is equipped with a 6-DoF Airbot Play robotic arm.
-The arm’s end effector is fitted with a single DoF parallel gripper [11], as shown in Figure 3.
+We use Isaac Gym to collect data from 4,096 agents for whole-body policy training.
+The policy is updated every 50 simulation steps and trained on an NVIDIA RTX 3090Ti, requiring about 8 hours for 10,000 iterations.
+As shown in Figure 3, the controller is deployed on a 12-DoF Uniree Go2 quadruped robot with a 6-DoF Airbot Play robotic arm.
+The arm’s end-effector is fitted with a single DoF parallel gripper [Zhaxizhuoma2025FastUMI].
 The robot total weights $19.9\text{\,}\mathrm{kg}$.
 Both the arm and the quadruped robot are powered by a battery on Go2.
-Inference for the whole-body policy runs onboard the Go2’s NVIDIA Jetson Orin NX, while the DP in automated mode is inferred on a PC equipped with an NVIDIA RTX 3090Ti, with the target trajectory sent via Ethernet.
-In teleoperation mode, a handheld gripper transmits the trajectory and gripper commands over the same link.
-During deployment, it is necessary to obtain the target pose and the current TCP pose in the world coordinate system.
-We install a Livox Mid-360 Lidar on the robot’s body and deploy an odometry [34] to build the world coordinate system, with the initial position of the Lidar serving as the origin.
-During system operation, the pose of the TCP is calculated through forward kinematics (FK) and transformed into the world coordinate system.
-The target pose is processed similarly.
-IV-B Ablation Studies
-We demonstrate the necessity and effectiveness of the proposed method through a series of ablation experiments conducted in simulations.
-For experiments with the same setup, we executed 20 trials with different seeds to record the data.
-Policy network wo TVP: Using the single MLP (sMLP) network-based Actor from [16] as a baseline, we validate the necessity of our TVP policy network for control.
-In the absence of external perception, quadruped robot movement relies on proprioceptive information to sense the environment.
-Our training results indicate that, when the component that estimates state using historical proprioceptive information is removed and a single MLP maps the history proprioceptive observation to actions, the task-tracking reward for loco-manipulation tasks rises more slowly during training and converges at a lower level, as shown in Figure 4.
-Evaluation further shows that this policy causes larger joint torque and position fluctuations, resulting in more pronounced jitter in joint motions, as shown in Figure 5(a) and (b).
-Figure 4: The trajectory tracking reward (left) and total reward (right) curve during the training of three policy networks: TVP (Ours), TVP without NAE, and single MLP.
-TVP policy network wo NAE: When the future target trajectory is unobservable (teleoperation), we carried out an ablation experiment to assess the performance conferred by the NAE, which predicts forthcoming trajectories from past trajectory data.
-In this setting, only the historical and current trajectories are encoded, and the resulting latent embedding is fed directly into the low-level network to generate action outputs.
-Because the policy lacks information about the trajectory it subsequently track, its tracking performance deteriorates: during training, the tracking reward remains lower than that achieved by the policy equipped with the NAE-based prediction mechanism, as illustrated in Figure 4.
-Figure 5(c) shows the MSE of the future trajectory predicted by the NAE decoder and its ground truth.
-The reconstruction error is represented by the position and the first two rows of pose matrix for four frames and plotted as a heatmap.
-The results demonstrate that the reconstruction error is minimal, indicating that the latent embedding of the NAE contains information essential for predicting the future trajectory.
-Figure 5: Policy network ablation analysis: (a) Schematic of the ‘unplug charger’ and ‘pushing’ action sequences in simulation.
-(b) Arm joint torque and position curves for TVP and single MLP networks during the tasks, with sMLP showing more jitter.
-(c) MSE heatmap of future trajectory reconstruction by TVP’s NAE for both tasks.
-Table III shows the policy evaluation for seven tasks.
-The position and orientation errors of the end-effector’s trajectory tracking were recorded after 800 time steps.
-Position error is measured using Euclidean distance (cm), while orientation error refers to the minimum rotational angle between the end-effector and the target trajectory’s orientation (radian).
+We implement asynchronous communication for our system through multi-thread.
+The whole-body policy runs onboard the Go2’s NVIDIA Jetson Orin NX at $50\text{\,}\mathrm{H}\mathrm{z}$.
+Inference of DP and the handheld gripper run on a PC with an NVIDIA RTX 3090 Ti that transmits the target trajectory via Ethernet.
+During deployment, it is necessary to obtain the target and end-effector poses in the world frame.
+A Livox Mid-360 LiDAR is mounted on Go2, and an odometry [Chen2023Direct] is used to build the world frame, with the LiDAR’s initial position as the origin.
+During inference, the poses are calculated through forward kinematics and transformed into the world frame.
+Figure 3: A multi-computer, multi-thread system deployment architecture.
+The host computer captures the handheld gripper’s trajectory ($100\text{\,}\mathrm{H}\mathrm{z}$) or runs the DP ($10\text{\,}\mathrm{H}\mathrm{z}$).
+The trajectory is sent to the lower computer via Ethernet.
+The Jetson Orin NX, mounted on the robot, handles multi-thread tasks, including receiving robot states ($500\text{\,}\mathrm{H}\mathrm{z}$), whole-body policy inference ($50\text{\,}\mathrm{H}\mathrm{z}$), and sending control commands ($200\text{\,}\mathrm{H}\mathrm{z}$).
+IV-B Simulation Studies
+We demonstrate the necessity and effectiveness of the proposed method through simulation experiments. For experiments with the same setup, we executed 20 trials with different seeds.
+Policy network wo TVP: Using a single MLP (sMLP) actor from the state-of-the-art work [Huy2024UMIonLegs] as a representative baseline, which has demonstrated good performance on pushing and cup-picking tasks, we highlight the necessity of our TVP policy network for control.
+This baseline policy does not predict the base velocity.
+The historical and current trajectories, the current proprioceptive state and the last action are fed into a single MLP network to generate the joint action.
+Our results show that the baseline policy reward exhibits slower growth and achieves a lower level of convergence, as shown in Figure 4.
+Evaluation further shows that this baseline policy causes larger joint torque and position fluctuations, leading to more noticeable joint jitter, as shown in Figure 5(a) and (b).
+To further quantify the jitter differences, we introduce a mean squared derivative (MSD) of arm joint torques over a time interval, $MSD=\frac{1}{(t_{1}-t_{0})/dt}{\textstyle\sum_{t=t_{0}}^{t=t_{1}-dt}(\tau_{t+dt}-\tau_{t})^{2}}$.
+In our case, the time interval is from $t_{0}=5s$ to $t_{1}=10s$.
+The results of ‘unplug charger’ in Table III show that the TVP can achieve smoother torque control, as it has the lower MSD than sMLP.
+Figure 4: The trajectory tracking reward (left) and total reward (right) curve during the training of four policy networks: TVP (Ours), TVP wo NAE, single MLP, and DeepWBC.
+TABLE III: MSD of six joint torques.
+Method Joint 1 Joint 2 Joint 3 Joint 4 Joint 5 Joint 6
+TVP 0.0254 0.2660 0.2144 0.0021 0.0026 0.0029
+sMLP 0.0515 0.2867 0.2500 0.0098 0.0038 0.0039
+TVP policy network wo NAE: In teleoperation, where the future trajectory is unobservable, we conducted an ablation study to evaluate the NAE, which predicts future trajectories from past data.
+In this ablation setting, the historical proprioceptive state sequence is fed into the ME to produce the ME embeddings.
+The ME embeddings are used as inputs to the estimator, which predicts the body’s linear velocity.
+The predicted linear velocity, the ME embeddings, the historical and current trajectories, the current proprioceptive state, and the last action are fed into a low-level network to generate the joint action.
+As shown in Figure 4, without trajectory prediction, the policy lacks guidance on “where to go next”, leading to degraded tracking performance: the tracking reward remains lower than that of the NAE-based policy.
+Figure 5(c) shows the MSE between the NAE-predicted trajectory and the ground truth.
+The reconstruction errors of position and the first two pose-matrix rows over four frames are visualized as a heatmap.
+The results demonstrate that the reconstruction error is minimal, indicating that the NAE embeddings contain information essential for predicting the future trajectory.
+Figure 5: Policy network ablation analysis: (a) The ‘unplug charger’ and ‘pushing’ trajectories in simulation.
+(b) Arm joint torque and position curves for TVP and sMLP during the tasks, with sMLP showing more jitter.
+(c) MSE heatmap of future trajectory reconstruction by NAE for both tasks.
+Table IV shows the policy evaluation for seven tasks.
+The position and orientation errors of the end-effector’s trajectory tracking were recorded over 800 time steps.
+Position error is measured using Euclidean distance (cm), while orientation error refers to the minimum rotational angle (radian).
 The average error over all time steps was computed to assess overall tracking performance.
-Results show that the TVP, with its future trajectory prediction mechanism, outperforms other policies, while the single MLP-based policy performs the worst due to its limited information extraction and representation ability.
-Adaptive trajectory sampling: In this experiment, we set all robots to uniformly sample (Uni-sample) all task trajectories throughout the training process in order to assess the necessity of our adaptive trajectory sampling strategy.
+Results show that the TVP, with its future trajectory prediction mechanism, outperforms other policies.
+The sMLP policy performs the worst due to its limited information extraction and representation ability.
+The error sources could include the dynamics and noise of the command trajectory, the dynamics of the robots, and the policy network.
+Policy wo NAE, task-frame target and trajectory input: We also add a comparison with DeepWBC [Fu2023DeepWholeBody].
+DeepWBC is obtained as follows.
+We remove the NAE that predicts future trajectories and discard trajectories as the policy targets, relying instead on a single target point.
+The end-effector pose is defined in the body frame.
+Through the above setting, we can obtain a method similar to DeepWBC.
+The experimental results for each task are presented in Table IV.
+Because the target is defined in the world frame rather than the body frame, our policy is not sensitive to body motion.
+Because our method has continuous trajectories as reference targets and has a prediction mechanism, our policy’s tracking performance is better than that of DeepWBC.
+Adaptive trajectory sampling: To assess the necessity of adaptive trajectory sampling mechanism, we set all robots to uniformly sample (Uni-sample) task trajectories throughout the training.
 We record the reward curves for two typical tasks during training, as shown in Figure 6.
 The results show that uniform sampling cannot effectively balance task performance.
-Initially, the reward curves for tasks with different pose variations are close, but over time, the rewards diverge, with tasks requiring more pose variation converging to lower rewards due to insufficient attention (i.e., higher sampling rates).
+Initially, reward curves for tasks with different pose variations are close.
+Over time, the rewards diverge.
+Tasks with more pose variation converge to lower reward due to insufficient attention.
 In contrast, adaptive sampling maintains similar rewards across tasks, balancing tracking performance.
-Table III indicates that tasks with more frequent trajectory changes (e.g., ”open container”) see a significant decline in tracking accuracy.
+Table IV indicates that tasks with frequent trajectory changes (e.g., ”open container”) see a significant decline in tracking accuracy.
 Figure 6: The trajectory tracking reward for two tasks under adaptive sampling and uniformly sampling.
-TABLE III: Simulation ablation study: mean 6D tracking error for each task, and the variance of those mean errors across all tasks.
+Moreover, the TVP (Ours) and the TVP wo NAE have lower variances than the baselines (sMLP and DeepWBC).
+Although the TVP (Ours) has a similar variance to the TVP wo NAE, the TVP (Ours) has the lowest mean tracking errors across all tasks.
+Impact of prediction errors: To evaluate the impact of prediction errors on policy performance, we inject Gaussian noise with increasing magnitudes into both the position (1 cm, 2 cm, 10 cm) and orientation (0.05 rad, 0.1 rad, 0.5 rad) components of the predicted trajectory.
+The experimental results are reported in Table IV.
+As the prediction errors grow, the tracking errors of the loco-manipulation also rise, but the increase remains acceptable, and the tasks can still be completed successfully.
+In extreme cases, where the predictor outputs unreasonable future steps (e.g., Gaussian noise with a mean of 10 cm and 0.5 rad), the loco-manipulation still functions.
+The reason is that our method does not rely solely on the predicted trajectory but also on the historical ground-truth trajectory, which provides good robustness against prediction errors.
+TABLE IV: Simulation ablation study: mean 6D tracking errors for each task, and the variance of those mean errors across all tasks.
 Metrics: Euclidean distance / rotation-matrix angular difference.
@@
 TVP (Ours) 1.32/0.08 1.08/0.07 0.87/0.11 0.66/0.05 0.74/0.06 0.94/0.09 1.07/0.15 0.050/0.0012
-TVP wo NAE 1.78/0.11 1.49/0.07 1.30/0.13 1.26/0.07 1.43/0.05 1.31/0.10 1.78/0.22 -
-sMLP (baseline) 2.23/0.12 2.02/0.14 2.31/0.37 2.24/0.14 2.20/0.13 2.17/0.22 3.13/0.54 -
+TVP wo NAE 1.78/0.11 1.49/0.07 1.30/0.13 1.26/0.07 1.43/0.05 1.31/0.10 1.78/0.22 0.049/0.0032
+sMLP [Huy2024UMIonLegs] 2.23/0.12 2.02/0.14 2.31/0.37 2.24/0.14 2.20/0.13 2.17/0.22 3.13/0.54 0.133/0.0256
+DeepWBC [Fu2023DeepWholeBody] 8.32/0.51 8.05/0.21 6.83/0.49 7.35/0.21 7.09/0.23 9.11/0.30 7.22/0.61 0.663/0.0279
+Noise ($N(1,0.2^{2})/N(0.05,0.01^{2})$) 1.43/0.10 1.23/0.08 0.92/0.11 0.79/0.05 0.91/0.06 1.08/0.09 1.26/0.16 0.053/0.0013
+Noise ($N(2,0.4^{2})/N(0.1,0.02^{2})$) 1.59/0.11 1.32/0.08 0.99/0.12 0.88/0.05 1.07/0.06 1.22/0.10 1.33/0.17 0.058/0.0016
+Noise ($N(10,2^{2})/N(0.5,0.1^{2})$) 2.19/0.13 1.51/0.11 1.28/0.15 1.50/0.07 1.40/0.09 1.39/0.11 1.79/0.25 0.098/0.0035
 Uni-sample 1.30/0.07 1.11/0.08 0.96/0.13 0.79/0.06 1.69/0.10 1.18/0.11 1.41/0.21 0.088/0.0026
 IV-C Real-world Experiments
-We further validate the feasibility and effectiveness of our method through real-world experiments.
-The experiments include both teleoperation via a handheld gripper and DP-based inference execution.
-All tasks were accomplished using a unified trained whole-body policy.
-Our approach demonstrates excellent performance in handling targets across all scenarios, achieving zero-shot transfer from simulation to the real world.
+We further validate the feasibility and effectiveness of our method through real-world experiments, including teleoperation and DP inference.
+All tasks are executed with a unified whole-body policy, showing strong performance across tasks and enabling zero-shot transfer from simulation to reality.
 Please refer to our videos in the supplemental materials.
-Handheld Gripper Teleoperation: Figure 1 I–V showcase several real-world demonstrations of whole-body loco-manipulation using the teleoperation, such as cleaning up paper scraps, picking up trash, storing toys, unplugging chargers, and removing bolts to open doors.
+Handheld gripper teleoperation: Figure 1 I–V showcase real-world demonstrations of whole-body loco-manipulation via teleoperation, including cleaning papers, picking up trash, storing a toy, unplugging a charger, and removing a bolt to open a door.
 Figure 1 III demonstrates a long-range task.
-It can be observed that when performing ground-level tasks, such as picking up trash, the quadruped robot assists the arm by leaning downward, enabling the end-effector to reach lower positions.
-Conversely, tasks well above body height require lifting the body.
-When the target position lies well beyond the current workspace, the quadruped robot exploits its locomotion capability to reposition itself, thereby reaching the distant site.
+For ground-level tasks, such as picking up trash, the quadruped robot assists the arm by leaning downward, enabling the end-effector to reach lower positions.
+Tasks well above body height require lifting the body.
+When the target position lies beyond the current workspace, the quadruped robot leverages its locomotion to reposition itself.
 These behaviors demonstrate exceptional whole-body coordination capabilities.
-We also evaluated the tasks for training and recorded the tracking performance in Table IV.
-The results show that the real-world deployment achieves tracking performance comparable to that of the simulation.
-TABLE IV: Real-world experiment results. (cm/radian/%)
+Two tasks in Table V show that the real-world deployment achieves tracking performance comparable to that of the simulation.
+TABLE V: Real-world experiment results. (cm / radian / %)
 EE pos err EE rot err Success Rate
@@
 open container 1.24 0.036 85
-Inference through DP: As shown in Figure 1 VI–VII, we demonstrate the method’s ability to autonomously execute tasks through a pre-trained DP.
-Specifically, we input the information from a GoPro camera mounted on the robotic arm’s wrist, along with the arm’s states, into the DP.
-The DP then autonomously generates the trajectory of the target object.
+Inference through DP: As shown in Figure 1 VI–VII, we demonstrate the ability to autonomously execute tasks through a pre-trained DP.
+Specifically, we input the information from a GoPro camera mounted on the arm’s wrist, along with the arm’s states, into the DP.
+The DP then generates the trajectory to the target object.
 The RL whole-body policy is then used to control the joints for execution.
 Its performance is mainly limited by the results of DP inference.
-The experimental results indicate that our method can leverage the DP to plan and autonomously perform multi-task operations, as shown in Table IV.
+The experimental results indicate that our method can leverage the DP to plan and autonomously perform multi-task operations, as shown in Table V.
 V Conclusion
 We proposed MLM, a reinforcement learning framework that leveraged both real-world and simulation data to address multi-task whole-body loco-manipulation control for quadruped robots.
-To balance performance while learning multiple task trajectories, we integrate a data library with an adaptive sampling mechanism.
-To address the scenario where only historical trajectories can be observed during teleoperation, we propose a Trajectory-Velocity Prediction policy network to predict future trajectories, thereby improving policy execution performance.
+To balance performance across multiple tasks, we integrate a data library with an adaptive sampling mechanism.
+For teleoperation where only historical trajectories are observable, we introduce a Trajectory-Velocity Prediction policy network to predict future trajectories, enhancing execution performance.
 Velocity estimation supports task execution across different spatial ranges.
-Finally, simulations and real-world experiments on Go2 and Airbot Play robot highlight the good performance of the proposed method.
-In fact, our approach can be extended to enable quadruped robots to perform more tasks in new environments, fully utilizing their locomotion and manipulation capabilities.
-This depends on higher-level decision-making and planning.
-Future integration of vision-language models and navigation could address this limitation.
+Simulations and real-world experiments on Go2 and Airbot Play also demonstrate the effectiveness of the proposed method.
+Looking ahead, our method can be extended to more tasks in new environments via higher-level decision-making (e.g., vision–language models).
+Moreover, handling highly dynamic tasks may benefit from faster sensor update rates and better actuation response capabilities.